OpenStudy (anonymous):

How do i differentiate this and find the gradient at point x=2

OpenStudy (anonymous):

\[y = \frac{4}{3x-4}\]

OpenStudy (anonymous):

@Callisto @precal @SmoothMath

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

\[(\frac{ f }{ g })^ '=\frac{ f^'g-fg^' }{ g^2 }\]

OpenStudy (anonymous):

qutient rule

OpenStudy (anonymous):

\[f=4,g=3x-4\]

OpenStudy (anonymous):

i dont understnd

OpenStudy (anonymous):

\[f^'=0,g^'=3\] so just plug in

OpenStudy (anonymous):

@satellite73 @Pallavi06 @SmoothMath @experimentX

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=\frac{ 0(3x-4) -3(4)}{ (3x-4)^2 }=\frac{ -12 }{ (3x-4)^2 }\]

OpenStudy (anonymous):

\[\frac{ -12 }{ (3(2)-4)^2 }=-3\]

OpenStudy (anonymous):

aren't there any easy methd?

OpenStudy (anonymous):

no........this is the method..........

OpenStudy (anonymous):

\[4(3x-4)^{-2}\] how do i do this now

OpenStudy (anonymous):

u have to differentiate it??

OpenStudy (anonymous):

yh

OpenStudy (anonymous):

wht is chain rule

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=\frac{dz }{ dx }\frac{ dy }{ dz }\]

OpenStudy (anonymous):

\[\frac{ 4 }{ 3x-4 }=4(3x-4)^{-1}\] let\[z=3x-4\] \[\frac{dz}{dt}=3\] \[y=4z^{-1}\] \[\frac{ dy }{ dz}=-4z^{-2}\]

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=?\]

OpenStudy (anonymous):

\[3(-4z^{-2})=-12(3x-4)^{-2}\]